OpenStudy (sugarmochi):

how does simplifying and evaluating sin^2 20˚ + sin^2 70˚ equal to one?

jhonyy9 (jhonyy9):

sin^2 20° = (0,34)^2 sin^2 70° = 1 - cos^2 70° = 1 - (0,34)^2 in this way than we rewrite sin^2 20° + sin^2 70° = sin^2 20° +1 - cos^2 70° = = (0,34)^2 +1 - (0,34)^2 = 1 hope helped

jhonyy9 (jhonyy9):

@ganeshie8 your opinion please @TheSmartOne ty.

OpenStudy (pythagoras123):

This can be proven algebraically without finding the exact value of either trigonometric function. \[\sin^220^0 + \sin^270^0 = \sin^220^0-\cos^270^0+1\] \[=(\sin20^0-\cos70^0)(\sin20^0 + \cos20^0)+1\] Let sin20 be x. We can construct a triangle to find the relation between sin20 and cos70:  Since cosx=adjacent / hypotenuse, cos70 = x / 1 = x = sin20. Hence our expression simplifies to \[0(\sin20^0+\cos20^0)+1\] \[=1\] (answer)

OpenStudy (unklerhaukus):

\[\sin(\theta)=\cos(90°-\theta)\]\[\sin^2(\theta)+\cos^2(\theta)=1\]

OpenStudy (divyankasc):

sin(θ)=cos(90°−θ) sin2(θ)+cos2(θ)=1